# Fishers Harvest Parallel Unlearning in Inherited Model Networks

### v1

@@ 6.1. General Label Unlearning Analysis
Figure 9. Unlearning layer analysis #$C_{f}$=15.
6.2. Merged Label Unlearning Analysis
This section assesses the applicability of the merging FIM function to models with various distributions of unlearned labels. We analyze the unlearning performance of different types of models across various datasets within the FL and DDPL frameworks. As demonstrated in Tables 6 and 7, with more experiments provided in Appendix .2.2, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method. We further validate that in the FIM, the impact of data on model parameters remains relatively stable.
We also analyze whether our proposed merging FIM function can handle the merged label unlearning task within the FL framework. As illustrated in Figure 7, in the AlexNet and DenseNet161 models, our method shows an average performance improvement of 88.54%, compared to using FIM directly. In the ResNet18 model, our method’s performance closely matches that of FIM. Additionally, our proposed merged label unlearning is, on average, 2.3 times faster than general label unlearning. This further validates the effiency of our method.
6.3. Inherited Depth of Model Analysis
We further examine whether the performance of FIUn is influenced by the increasing depth of inheritance within the FL and IL frameworks. All experiments are conducted with the same hyperparameters.
@@ 6.3. Inherited Depth of Model Analysis
  • Federated Learning. In the FL framework, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments. Each model trains using one segment of the data, adopting a binary tree structure as depicted in Figure 6. We evaluate whether our method can effectively unlearn specified labels as the depth of the binary tree increases.
@@
  • Incremental Learning. In the IL framework, as demonstrated in Figure 6, the starting model $w_{g}$ contains data from 170 classes. With each incremental learning step, one class of label data is added to assess whether our method can effectively unlearn specified labels.
As shown in Figure 8 (i.e., single-class unlearning, see Appendix .4), we observe that under both frameworks, various datasets, and models with different unlearned labels, the performance of the proposed algorithm remains consistent regardless of the increased inherited depth. This consistency across both frameworks and models supports the notion that the impact of data on model parameters remains stable at fixed positions. Such stability ensures that our method maintains high performance even as the depth increases.
6.4. Analysis of Parameter Layer Selection in Models
@@
We analyze how the number of model layers used to calculate the FIM impacts the performance of our algorithm. We conduct experiments within both FL and IL frameworks, with 15 label categories selected as unlearned labels. The DAG diagrams for these frameworks are illustrated in Figures 6 and 6. We calculate the FIM using the last layer, the last four layers, the last seven layers, and all layers to assess unlearning performance.
As depicted in Figure 9, we observe that an increase in the number of layers used leads to diminishing unlearning performance across various datasets and models. This decline is attributable to the increase in the layer parameters, which inversely affects the $AD_{r}$ and $AD_{f}$ metrics. Moreover, a larger number of layers prolongs the time needed to compute the FIM. This finding underscores the effectiveness of our method, which relies on calculations from only the last layer.
7. Related Work
Although FL (Li et al., 2020b), DDPL (Li et al., 2020c), IL (Van de Ven et al., 2022), and TL (Weiss et al., 2016) have each made significant strides in addressing specific challenges, such as data privacy protection (Li et al., 2021), model training efficiency (Dean et al., 2012), knowledge retention and unlearning (Han et al., 2023), and cross-domain knowledge transfer (Rajapaksha et al., 2023), systematic exploration that combines these learning frameworks remains sparse. Most studies have merely attempted simple combinations of a few methods (Dai and Meng, 2023; Yoon et al., 2021; Dai et al., 2019; Miao et al., 2023) without touching the core of their commonality–model inheritance. This characteristic plays a crucial role in environments where unlearning is conducted on a large-scale. As data and models grow, the computational overhead increases exponentially with the depth of inheritance, while time consumption grows linearly with the depth, even when unlearning can be executed simultaneously at the same depth starting from the unlearning origin. This holds true regardless of the existing unlearning methods utilized, which are respectively discussed in the following.
Federated Unlearning. FedEraser (Liu et al., 2020) used historical parameter updates from the central server to reconstruct the deleted model. Wu et al. (Wu et al., 2022a) subtracted historical updates and used knowledge distillation to restore model performance without client data. FRU (Yuan et al., 2023) recalled and calibrated historical updates to eliminate user contributions and accelerate federated recommendation reconstruction. These methods primarily relied on recalling and calibrating historical parameter updates of the model to achieve deletion or re-training. Halimi et al. (Halimi et al., 2022) reversed the learning process by maximizing local empirical loss and used PGD to perform deletion on the client side, while Wu et al. (Wu et al., 2022b) used class, client, and sample decoupling learning combined with Stochastic Gradient Ascent (SGA) and Elastic Weight Consolidation (EWC) to achieve deletion. These methods achieved deletion by optimizing the model’s loss function or gradients.
Liu et al. (Liu et al., 2022) utilized first-order Taylor expansion and a diagonal empirical Fisher information matrix for accelerated re-training. Wang et al. (Wang et al., 2022) applied CNN channel pruning to remove specific category information, while FFMU (Che et al., 2023) employed nonlinear function analysis to enhance model deletion processes. These approaches streamline deletion or re-training by altering the model structure or employing targeted techniques. Additionally, KNOT (Su and Li, 2023) optimized re-training by introducing clustering aggregation and framing the client clustering issue as a dictionary minimization problem.
Distributed Data-Parallel Unlearning. In DDPL knowledge unlearning, handling the unlearning process is restricted to the data shard on the specific device, followed by model aggregation and iterative training. If unlearning is required from a specific shard, only the model trained on that shard is re-trained. Approximate machine unlearning directly modifies trained model parameters, employing methods such as 1) Gradient-based approaches (Fan et al., 2023; Golatkar et al., 2020; Neel et al., 2021), which update model gradients to gradually diminish the data’s influence, and 2) Data perturbation (Tarun et al., 2023; Goel et al., 2024), which introduces noise to reduce specific data points’ impact on the model. These methods facilitate effective knowledge unlearning in DDPL.
Incremental Unlearning. Research on unlearning on IL is still nascent, yet the Projected-Gradient Unlearning method (PGU) (Hoang et al., 2024) shows promise. PGU computes the gradients for the unlearning dataset and projects these onto the Core Gradient Space (CGS) derived from the training set. It then subtracts this projected component from the original gradients, positioning the residuals in a space orthogonal to the CGS, and updates the model parameters with these adjusted gradients. Effective for one-time data deletions and batch-wise unlearning, PGU is well-suited for ongoing incremental learning processes.
Transfer Unlearning. Research on unlearning specific target task knowledge has been relatively scarce in TL. To address this challenge, (Sepahvand et al., 2024) proposed fine-tuning pre-trained models through data selection. Subsequently, the entire network was fine-tuned on the target task using gradient descent.
What Sets Our Solution Apart. The above-mentioned existing studies on unlearning methods typically endure high computational complexity, significant resource demands, and implementation difficulties. These methods are time-consuming and resource-intensive, especially when deployed in large-scale environments, as they require accurate tracking and sequential updating of numerous model parameters. By contrast, the proposed FIUn method fully parallelizes unlearning, significantly enhancing computational efficiency, resource utilization, and adaptability. It effectively addresses the challenges faced by existing unlearning methods in inheritance-oriented model networks.
8. Conclusion
@@
In this paper, we developed a new parallel unlearning framework for models with dependence and inheritance. To do this, we introduced the UMIG, which can effectively abstract model inheritance, e.g., in FL, DDPL, IL, and TL, using a DAG topology. We also developed the FIUn method and the merging FIM function, which can quickly locate impacted models and conduct parallel unlearning through the UMIG. Experimental results on various models and datasets showed that our method achieves complete unlearning for single-class labels, while maintaining an average accuracy of 94.53% for retained labels. For multi-class labels, the unlearning accuracy is only 1.07%, with retained label accuracy at 84.77%. Our method is 99% faster than existing unlearning methods, efficiently handling model updates regardless of inheritance depth.

### v4

@@ VII-A General Label Unlearning Analysis
Figure 11: Unlearning layer analysis $C_{f}$=15 for all layers [foster2024fast, golatkar2020eternal], last 1 layer and last 7 layers.
TABLE IX: Large-scale Federated Unlearning Performance of Language and Graph Models on Yahoo! Answers
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
Bert-Base-Cased
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 105.63 215.14 213.75
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.99 0.00 0.00 1.56 2.84 2.79
2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 93.15 186.43 190.32
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 15.13 0.00 10.21 1.51 2.75 2.81
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 70.31 143.53 145.17
$AD_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 26.73 0.00 6.41 1.90 2.91 3.14
GCNN
1 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.69 99.99 97.64 91.54 68.69 95.67 33.95 67.31 101.64
$AD_{f}\downarrow$ 98.81 99.99 94.64 0.00 0.00 0.00 0.00 0.00 9.52 0.42 0.77 0.76
2 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.31 63.89 64.29 71.64 28.60 57.14 86.41
$AD_{f}\downarrow$ 99.11 99.99 95.60 0.00 0.00 0.00 0.00 0.00 0.00 0.44 0.87 0.86
4 $AD_{r}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.34 93.17 63.07 97.41 21.63 42.60 64.31
$AD_{f}\downarrow$ 99.07 99.99 95.67 0.00 0.00 0.00 0.00 0.00 25.37 0.34 0.79 0.84
(a) Throughput, FL, TinyImageNet, ResNet18
(b) Latency, FL, TinyImageNet, ResNet18
(c) Throughput, FL, TinyImageNet, DenseNet161
(d) Latency, FL, TinyImageNet, DenseNet161
(e) Throughput, FL, Yahoo, Bert
(f) Latency, FL, Yahoo, Bert
(g) Throughput, FL, Yahoo, GCNN
(h) Latency, FL, Yahoo, GCNN
Figure 12: Scalability results of the FIUn method in FL. Settings: unlearning requests from 1,000 to 10,000 (step 1,000), 8,000 concurrent requests, DAG depth 3, models include TinyImageNet (ResNet18, DenseNet161) and Yahoo (Bert, GCNN). Line charts show throughput vs. transactions; bar charts show latency.
VII-B Merged Label Unlearning Analysis
This experiment assesses the applicability of the MFIM function to models with various distributions of unlearned labels. We analyze the unlearning effectiveness of various models across CIFAR100 DDPL. As demonstrated in Table VII, the average $AD_{r}$ is 89.72%, and the average $AD_{f}$ is 5.62%, confirming the unlearning effectiveness of our method. We evaluate unlearning on TinyImageNet under FL. Table VIII shows that FIUn achieves complete forgetting of target classes while keeping retained-class accuracy high (92.4%), demonstrating strong stability and generalization via the MFIM function. FIUn also greatly reduces cumulative unlearning time compared to retraining and maintains consistently low unlearning time across label distributions. Additional experiments are provided in Appendix B.
We evaluate the MFIM function’s ability to manage merged label unlearning tasks within FL. As shown in Fig. 8, our method achieves an average 88.54% performance improvement in AlexNet and DenseNet161 models over direct FIM use. In the ResNet18 model, its unlearning effectiveness is comparable to that of FIM. This is because MFIM reduces cumulative errors (caused by the manual threshold rule) by merging FIMs from all relevant roots into a single calculation, allowing parameter adjustments to be made in one step instead of applying the threshold rule to each root [foster2024fast, golatkar2020eternal], leading to more accurate and stable unlearning outcomes.
In experiments with BERT-base-cased and GCNN on the Yahoo! Answers dataset within FL (Table IX), we tested 50 clients for multi-label unlearning. The $AD_{f}$ metric showed small increases in some cases (15.13% and 26.73%), but overall remained low. For GCNN, under $C_{f}$=2 and 4, $AD_{r}$ exceeded 63%, though FIUn slightly reduced retained accuracy due to the large, unoptimized FIM hyperparameter space. Nevertheless, FIUn was highly time-efficient, significantly faster than re-training, and achieved effective unlearning with only minor accuracy trade-offs.
Takeaway–capable of multi-root unlearning. With up to 88.54% improvement, the MFIM function enables FIUn to consistently manage merged unlearning across different tasks and models, a capability that standard FIM approaches cannot achieve due to their inability to aggregate multiple FIMs into a single and unified FIM that encapsulates all root updates.
VII-C Inherited Depth of Model Analysis
We assess FIUn’s performance impact as inheritance depth increases in FL and IL, using consistent hyperparameters across all experiments.
  • $\bullet$
@@ VII-C Inherited Depth of Model Analysis
    Federated Learning. (Cyan in Fig. 5) In FL, the CIFAR100 and TinyImageNet datasets are randomly divided into five segments. Each model trains using one segment of the data that adopts a binary tree structure. We evaluate whether our method can effectively unlearn specified labels as the depth of the binary tree increases.
  • $\bullet$
@@
    Incremental Learning. (Blue in Fig. 5) In IL, the starting model $w_{g}$ includes data from 170 classes. With each learning step, a new class is added to evaluate the effectiveness of our method in unlearning specific labels.
Figs. 9 and 10 evaluate inheritance depth under FL and IL for $C_{f}\in{1,10}$. Overall, our method remains stable across depths; in FL, the unlearning accuracy $\bigtriangleup_{acc}$ stays near 100% even at large depths. In IL, especially TinyImageNet + DenseNet161, $\bigtriangleup_{acc}$ declines and fluctuates as depth grows, likely because DenseNet161’s dense connectivity accumulates class-specific features that are harder to erase, and TinyImageNet’s strong inter-class overlap amplifies category and parameter drift; depth-dependent adaptation and data stochasticity further contribute to variability.
Takeaway–perfectly matching inheritance networks. FIUn exhibits the Hyper-Distance property, effectively unlearning across up to 50 levels of inheritance depth in various models and datasets, enabling simultaneous updates of all inherited models during unlearning tasks.
VII-D Parameter Layer Selection in Models
@@
We evaluate the impact of the number of model layers used for FIM calculation on algorithm performance. Experiments were conducted in FL and IL frameworks with 15 label categories selected for unlearning. Fig. 5 illustrates the DAG for these frameworks with Yellow and Blue models ($w_{g}$, $w_{a}$, $w_{b}$). FIM calculations were conducted on the last layer, last seven layers, and all layers using 10 random parameter sets.
We apply the methods from [foster2024fast, golatkar2020eternal] (using all layers) for unlearning in inherited models. As shown in Fig. 11, computing FIM on all layers leads to high variance in $\bigtriangleup_{acc}$. In contrast, the “Last 1” and “Last 7” settings show much smaller variance and achieve over 100% better unlearning. This suggests that focusing on partial layers—especially the last layer—yields more stable and consistent results. The reason is that early layers capture general features, while later layers, particularly the last, encode task-specific features. Thus, targeting the last layer removes task-specific information more efficiently, which fits well with our MFIM design.
Takeaway–focusing on the last layer is the best for efficiency and unlearning effectiveness. FIUn achieves more efficient and effective unlearning by strategically performing FIM calculations on the weights of the last layer of models that the calculation of the MFIM function can benefit from.
VII-E Unlearning Speed Analysis
We conduct experiments on the inherited model $w_{b}$ using AlexNet and ResNet models, as shown in Fig. 5, while the unlearning speed is presented in Fig. 13. The results show that the unlearning time progressively increases with the number of unlearning layers, the parameters for AlexNet are 409600, 18833,764, and 20495268 for 1, 7, and all layers, respectively, while for ResNet18, they are 51200, 4903524, and 11227812, respectively. The results confirm that our proposed FIUn method with the last layer effectively enhances the efficiency of the unlearning process.
Figure 13: Unlearning speed of different layers of models. FIUn speed is based on the FL framework with CIFAR100.
VII-F Scalability Experiments
To validate the scalability of FIUn, we conduct parallel experiments with the following key configurations:
  • Unlearned data volume: Requests range from 1,000 to 10,000, in increments of 1,000.
  • Concurrent requests: A concurrency level of 8,000 is maintained.
  • Distributed participants: Experiments use a DAG depth of 3, with scalability tested on federated setups across models such as TinyImageNet (ResNet18, DenseNet161) and Yahoo (Bert, GCNN).
Fig. 12 reports FL scalability with FIUn, measuring throughput and latency on TinyImageNet (ResNet18, DenseNet161) and Yahoo (BERT, GCNN). FIUn shows near-linear throughput growth and stable latency as transaction volume increases, reflecting its efficient parallel unlearning; the setup is reproducible and indicates real-world applicability. Additional experiments appear in Appendix C.
VIII Conclusion
@@
We developed a novel parallel unlearning framework for models with dependence and inheritance, utilizing a chronological DAG to abstract and visualize model inheritance in FL, DDPL, IL, and TL. The FIUn method and MFIM function efficiently identified affected models and enabled parallel unlearning through the DAG. Experiments across various models and datasets demonstrated that FIUn achieved complete unlearning for single-class labels while maintaining an average retained accuracy of 94.53%. For multi-class labels, unlearning accuracy was just 1.07%, with retained accuracy at 84.77%. Our method is 99% faster than existing ones and handles model updates efficiently at all inheritance depths.
TABLE X: Federated Unlearning Performance on CIFAR100
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet
1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 97.13 96.41 97.12 29.70 59.06 89.62
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.13
2 $AD_{r}\uparrow$ 99.99 99.99 99.86 99.99 99.99 99.51 91.89 90.78 95.66 28.69 57.81 84.97
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.16 0.13
4 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 85.79 78.54 89.40 26.39 53.10 81.96
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.10 0.16 0.13
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 97.09 97.13 96.93 30.36 61.40 90.94
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.68 1.00 0.99
2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 96.66 98.10 97.44 28.30 57.39 84.06
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.69 1.03 1.09
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 97.17 99.45 99.19 26.39 52.69 81.39
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.75 1.09 1.06
TABLE XI: Incremental Unlearning Performance on CIFAR100
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet
1 $AD_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 98.97 94.39 94.98 22.95 43.83 63.45
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.39 0.39
2 $AD_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 92.93 71.58 69.14 21.35 41.58 62.10
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.27 0.26
4 $AD_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 88.77 59.62 58.81 20.52 41.12 60.69
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00 0.12 0.28 0.29
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.71 90.05 90.82 26.35 47.12 66.93
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.30 0.80 0.85
2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.49 89.02 87.64 25.37 45.93 65.89
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.30 0.84 0.85
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 84.71 83.42 84.54 24.83 43.23 65.38
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.31 0.86 0.87
TABLE XII: Transfer Unlearning Performance on TinyImageNet
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161
1 $AD_{r}\uparrow$ 86.33 91.24 89.34 97.75 97.52 96.75 71.49 69.82 66.73 987.94 1986.35 1985.16
${AD}_{f}\downarrow$ 94.50 94.50 96.70 0.00 0.00 0.00 0.00 0.00 0.00 2.41 4.15 4.34
4 $AD_{r}\uparrow$ 86.33 91.24 89.34 96.42 96.41 95.13 64.24 55.32 51.18 1001.34 1994.53 1986.61
${AD}_{f}\downarrow$ 89.12 91.23 91.84 0.00 0.00 0.00 0.00 0.00 0.00 2.34 4.71 4.26
6 $AD_{r}\uparrow$ 86.33 91.24 89.34 95.72 95.43 93.58 58.69 47.38 42.57 1018.14 2008.89 2009.31
${AD}_{f}\downarrow$ 86.57 89.25 90.28 0.00 0.00 0.00 0.00 0.00 0.00 2.16 4.64 4.46
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 98.07 96.74 142.15 243.75 241.61
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.42 0.75 0.78
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 98.38 97.38 143.46 244.34 245.74
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.42 0.74 0.75
6 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 98.82 98.91 98.98 147.27 245.35 246.41
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.56 0.89 0.74
TABLE XIII: Distributed Data-Parallel Unlearning Performance on TinyImageNet
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161
1 $AD_{r}\uparrow$ 99.99 87.16 85.21 99.99 89.39 87.24 91.82 86.63 83.58 980.53 988.42 991.75
${AD}_{f}\downarrow$ 99.99 92.75 90.41 0.00 0.00 0.00 0.00 0.00 0.00 4.34 2.61 2.13
4 $AD_{r}\uparrow$ 99.96 87.00 85.21 99.96 88.35 89.34 87.42 86.56 84.37 1004.63 1006.32 1005.332
${AD}_{f}\downarrow$ 99.96 86.97 88.85 0.00 0.00 0.00 0.00 0.00 0.00 4.31 2.05 2.35
6 $AD_{r}\uparrow$ 99.99 87.00 85.21 99.96 87.96 88.15 82.50 86.82 84.47 1024.25 1036.14 1026.74
${AD}_{f}\downarrow$ 99.99 89.35 86.08 0.00 0.00 0.00 0.00 0.00 0.00 4.16 2.24 2.21
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.19 99.99 99.99 112.52 117.34 114.52
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.56 1.67 1.47
4 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 99.99 109.32 108.74 110.53
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.32 3.54 1.74 1.37
6 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.96 99.96 99.96 99.38 99.99 99.99 107.78 108.32 107.17
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.53 1.84 1.64
Appendix A General Label Unlearning Analysis
Table X to Table XIII show that, on CIFAR100, FIUn achieves strong performance under federated unlearning and incremental unlearning frameworks. Similarly, on TinyImageNet, FIUn exhibits excellent performance under transfer unlearning and distributed data-parallel unlearning. Re-training on TinyImageNet takes a significantly longer time than CIFAR100. In comparison, FIUn maintains a high unlearning accuracy while keeping the unlearning time extremely short.
Appendix B Merged Label Unlearning Analysis
Tables XIV to XVII show that, in the federated unlearning setting on CIFAR100, our method achieves good performance on multi-label distributions. On TinyImageNet, under the distributed data-parallel unlearning setting, it demonstrates excellent performance on multi-label distributions. For multiple label distributions, our proposed MFIM function accelerates the unlearning speed of inherited models while maintaining high unlearning accuracy.
Appendix C More Datasets and Large Scale Analysis
Tables XV and XVI present the experimental results of unlearning using a ViT model on ImageNet and TinyImageNet datasets. Table XV shows that for ImageNet, when unlearning one class ($C_{f}=1$), the accuracy on retained data ($AD_{r}$) remains stable (73.8% $\rightarrow$ 73.64%) while the accuracy on unlearned data ($AD_{f}$) drops from 100% to 0%, with an unlearning time of 2.1s. For ten classes ($C_{f}=10$), $AD_{f}$ decreases from 89.04% to 0%, with $AD_{r}$ slightly reduces to 73.1% and an unlearning time of 4.33s. In the TinyImageNet experiment, as shown in Table XVI, we simulate a federated learning setup with 100 clients, each running the same ViT model. For both $C_{f}=1$ and $C_{f}=10$, $AD_{f}$ reaches 0% after unlearning, while $AD_{r}$ remains above 95%, with unlearning times ranging from 3.79s to 13.14s across clients.
TABLE XIV: Federated Unlearning Performance on CIFAR-100 with Multiple Label Distribution
Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
AlexNet
60% $AD_{r}\uparrow$ 99.99 99.99 99.81 98.80 99.99 99.99 69.42 81.68 75.05 59.26 30.13 29.70
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.08 0.09
40% $AD_{r}\uparrow$ 99.99 99.99 99.80 99.51 99.99 99.99 61.66 79.77 72.96 53.84 29.58 28.69
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00 0.11 0.09 0.19
20% $AD_{r}\uparrow$ 99.99 99.99 99.96 99.16 99.99 99.99 59.57 77.47 77.97 57.36 27.48 26.39
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.12 0.00 0.00 0.10 0.07 0.18
ResNet18
60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 93.15 98.76 97.81 67.31 33.18 30.36
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.60 0.32 0.29
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 93.78 99.04 98.98 52.33 28.13 28.30
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.59 0.34 0.30
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 88.54 98.79 98.62 56.61 23.37 26.39
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.58 0.28 0.30
TABLE XV: ImageNet Experiment Results for Unlearning with Pretrained ViT Model
Unlearning Label AD_r (%) AD_f (%) AD_r after Unlearning (%) AD_f after Unlearning (%) Unlearning Time (s)
One Class ($C_{f}=1$) 73.8 100 73.64 0 2.1
Ten Classes ($C_{f}=10$) 73.8 89.04 73.1 0 4.33
TABLE XVI: TinyImageNet Experiment Results for Unlearning in Federated Learning Setup with ViT Model
Unlearning Case Client Unlearning Time (s) $AD_{f}$ after Unlearning (%) $AD_{r}$ after Unlearning (%)
One Class ($C_{f}=1$) Client 1 3.79 0.00 95.63
Client 2 7.94 0.00 99.98
Client 3 12.45 0.00 99.63
Ten Classes ($C_{f}=10$) Client 1 3.84 0.00 99.99
Client 2 7.96 0.00 99.99
Client 3 13.14 0.00 99.99
TABLE XVII: Distributed Data-Parallel Unlearning Performance on TinyImageNet with Multiple Label Distribution
Model $\cap_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Unlearning Time (s)
Re-training FIUn
$w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$ $w_{g}$ $w_{a}$ $w_{b}$
DenseNet161
60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 76.86 89.47 82.53 994.32 998.46 1000.66
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 4.69 2.39 2.30
40% $AD_{r}\uparrow$ 99.99 99.99 99.50 99.99 99.99 99.99 74.03 89.47 83.18 1010.45 1011.14 1009.53
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 4.12 2.69 2.14
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 73.48 89.13 83.78 1021.34 1020.67 1022.64
${AD}_{f}\downarrow$ 99.99 99.99 99.91 0.00 0.00 0.00 0.00 0.00 0.00 4.98 2.64 2.59
ResNet18
60% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 95.16 95.92 98.86 123.74 126.43 149.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.16 1.65 1.68
40% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 94.43 95.92 98.27 128.43 127.64 155.32
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.97 1.46 1.59
20% $AD_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 94.71 95.92 98.48 129.35 126.43 154.83
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 3.89 1.54 1.38